The acceleration ‘a’ (in ms–2) of a body, starting from rest varies with time t (in s) following the equation a = 3t + 4
The velocity of the body at time t = 2s will be
(A) 10 ms–1 (B) 18 ms–1 (C) 14 ms–1 (D) 26 ms–1

Grade:

2 Answers

Arun
3 years ago
Given => 
a = 3t + 4
We can find the velocity by integrating both sides and by putting in the limits!
So, by integration of both sides,
we get,​
v=[3t2/2 + 4t]
As t=2, by putting in the value, we get!
v = 3(2)2/2 + 4(2)
v= 6+8
v=14ms-1
 
Vikas TU
3 years ago
a = 3t + 4
We can find the velocity by integrating both sides and by putting in the limits!
So, by integration of both sides,
we get,​
v=[3t2/2 + 4t]
As t=2, by putting in the value, we get!
v = 3(2)2/2 + 4(2)
v= 6+8
v=14ms-1